Two trains started fromt he same station at the same time and traveled in opposite directions. After traveling 10 hours, they were 1,400 miles apart. The rate of the fast train exceeded the rate if the slow train by 5 miles per hour. Find the rate of each train.
Can you please help me solve this equation? Thank you.

Answer by nerdybill(6959)

Two trains started fromt he same station at the same time and traveled in opposite directions. After traveling 10 hours, they were 1,400 miles apart. The rate of the fast train exceeded the rate if the slow train by 5 miles per hour. Find the rate of each train.
.
Let x = speed of slow train
then
x+5 = speed of fast train
.
Using the "distance formula":
d = rt
where
d is distance
r is speed or rate
t is time
.
10x + 10(x+5) = 1400
10x + 10x + 50 = 1400
20x + 50 = 1400
20x = 1350
x = 67.5 mph (slow train)
.
Fast train:
x+5 = 67.5+5 = 72.5 mph